# Assessing AI Rationality: The Random Guesser Test for Sequential Decision-Making Systems

### v1

@@ Abstract.
We propose a general approach to quantitatively assessing the risk and vulnerability of artificial intelligence (AI) systems to biased decisions. The guiding principle of the proposed approach is that any AI algorithm must outperform a random guesser. This may appear trivial, but empirical results from a simplistic sequential decision-making scenario involving roulette games show that sophisticated AI-based approaches often underperform the random guesser by a significant margin. We highlight that modern recommender systems may exhibit a similar tendency to favor overly low-risk options. We argue that this “random guesser test” can serve as a useful tool for evaluating the rationality of AI actions, and also points towards increasing exploration as a potential improvement to such systems.
@@ 3.1. Sequential Gambling
We use two different sets of winning probabilities. One is a fair roulette setting, where the expected values of all the bet types are the same, and a skewed roulette setting, where the zero bet has a positive expected value of \$1. In either case, the AI agent is unaware of the true nature of the wheel. If the AI agent is intelligent enough, however, it should be able to uncover these hidden rules behind the data. In other words, a rational and effective AI should be able to detect any real and exploitable patterns within the game.
@@
The skewed roulette setting is used to assess the impact of data non-stationalities. The importance of different stationalities in a similar scenario have also been discussed in Lin et al. (Lin et al., 2019). Most recent works also assume that the underlying data is stationary, which makes real-life applications questionable. In our scenario, we manifest this with a situation in which the roulette wheel is skewed temporarily such that the chance to win the zero bet doubles. Again, our agent is unaware that this change has occurred. This extra element allows us to test whether or not these agents are rationally able to explore and exploit this information.
@@ 4. Empirical Evaluation
This section presents the results of the empirical evaluation of the decision-making algorithms introduced in the previous section.111 The Python implementation will be made publicly available upon acceptance of the paper. We employed two metrics to evaluate utility. One is the rate of successful sessions under a finite session horizon $T$, where a session (i.e., a sequence of bet rounds) is considered successful when the balance is nonnegative at the completion of the final round $t=T$. The other metric we used is the number of rounds until bankruptcy, defined as the number of bet rounds until the balance becomes zero or below. For the latter, no finite session horizon is assumed, and hence, the metric can theoretically be positive infinite.

### v2

# Assessing AI Utility: The Random Guesser Test for Sequential Decision-Making Systems

@@ Abstract.
We propose a general approach to quantitatively assessing the risk and vulnerability of artificial intelligence (AI) systems to biased decisions. The guiding principle of the proposed approach is that any AI algorithm must outperform a random guesser. This may appear trivial, but empirical results from a simplistic sequential decision-making scenario involving roulette games show that sophisticated AI-based approaches often underperform the random guesser by a significant margin. We highlight that modern recommender systems may exhibit a similar tendency to favor overly low-risk options. We argue that this “random guesser test” can serve as a useful tool for evaluating the utility of AI actions, and also points towards increasing exploration as a potential improvement to such systems.
@@ 3.1. Sequential Gambling
We use two different sets of winning probabilities. One is a fair roulette setting, where the expected values of all the bet types are the same, and a skewed roulette setting, where the zero bet has a positive expected value of \$1. In either case, the AI agent is unaware of the true nature of the wheel. If the AI agent is intelligent enough, however, it should be able to uncover these hidden rules behind the data. In other words, an effective AI agent should be able to detect any real and exploitable patterns within the game.
@@
The skewed roulette setting is used to assess the impact of data non-stationalities. The importance of different stationalities in a similar scenario have also been discussed in Lin et al. (Lin et al., 2019). Most recent works also assume that the underlying data is stationary, which makes real-life applications questionable. In our scenario, we manifest this with a situation in which the roulette wheel is skewed temporarily such that the chance to win the zero bet doubles. Again, our agent is unaware that this change has occurred. This extra element allows us to test whether or not these agents are effectively able to explore and exploit this information.
@@ 4. Empirical Evaluation
This section presents the results of the empirical evaluation of the decision-making algorithms introduced in the previous section.111 A Python implementation is available at https://github.com/ShunIde/Python-Implementation-for-AIBS-2024-Submission We employed two metrics to evaluate utility. One is the rate of successful sessions under a finite session horizon $T$, where a session (i.e., a sequence of bet rounds) is considered successful when the balance is nonnegative at the completion of the final round $t=T$. The other metric we used is the number of rounds until bankruptcy, defined as the number of bet rounds until the balance becomes zero or below. For the latter, no finite session horizon is assumed, and hence, the metric can theoretically be positive infinite.